OpenStudy (anonymous):

A store can develop 225 photos every half hour. A. Complete the rate table. Number of Photos 225 Time (hours) 0.5 1 3 4 12 B. How many photos can the store develop in 12 hours?

OpenStudy (texaschic101):

225 every half hour 0.5 = 225 1 = 450 (per hr) 3 = 450(3) = 1350 4 = 450(4) = 1800 12 = 450(12) = 5400